In a family there are two cars. In a given week, the first car gets an average of 40 miles per gallon, and the second car gets 30 miles per gallon. The two cars combined drive a total of 2300 miles in that week, for a total gas consumption of 65 gallons. How many gallons were consumed by each of the two cars that week?

Answer by stanbon(75887)

In a family there are two cars. In a given week, the first car gets an average of 40 miles per gallon, and the second car gets 30 miles per gallon. The two cars combined drive a total of 2300 miles in that week, for a total gas consumption of 65 gallons. How many gallons were consumed by each of the two cars that week?
----------------------
Note: distance/mpg = gallons
---------------------------------
1st car DATA:
mpg = 40 ; distance = x miles ; gallons = x/40
-----------------------------------
2nd car DATA:
mph 30 ; distance = 2300-x ; gallons = (2300-x)/30
-------------------------------------
Equation:
gallons + gallons = 65
x/40 + (2300-x)/30 = 65
------------
Multiply thru by 120 to get:
3x + 4(2300-x) = 65*120
3x -4x + 4*2300 = 65*120
-x = -1400
x = 1400 miles
2300-x = 900 miles
----------------------------
x/40 = 35 gallons (amt. used by the 1st car)
----
900/30 = 30 gallons (amt. used by 2nd car)
==========================================
Cheers,
Stan H.